OpenStudy (laddiusmaximus):

@satellite73 find the equation of the line perpendicular to the line 2x+5y=12 and through the midpoint AB where A(-3,-4) and B (7,-8)

OpenStudy (anonymous):

fist you need the slope of \(2x+5y=12\) the slope of \(ax+by=c\) is \(-\frac{a}{b}\) so in your case the slope is \(-\frac{2}{5}\)

OpenStudy (anonymous):

slope of perpendicular line is the negative reciprocal, in your case \[m=\frac{5}{2}\]

OpenStudy (anonymous):

now we need the midpoint of \((-3,-4) (7,-8)\) take the average in each coordinate and get \[(\frac{7-3}{2},\frac{-4-8}{2})\] or \[(2,-6)\]

OpenStudy (whpalmer4):

@satellite73 since you won't accept messages from me: how did you apply the formatting in the first message in this thread?

OpenStudy (anonymous):

then point slope forumula \[y=y_1=m(x-x_1)\] with \[m=\frac{5}{2},x_1=2,y_1=-6\] and write \[y+6=\frac{5}{2}(x-2)\] solve for \(y\) if you like or whatever form you need

OpenStudy (anonymous):

@whpalmer4 latex see latex practice group for examples

OpenStudy (anonymous):

or right click on equation, select "show math as" then "latex"

OpenStudy (laddiusmaximus):

where did the 5/2 come from?

OpenStudy (anonymous):

the line \( 2x+5y=12\) has slope \(-\frac{2}{5}\) perpendicular line has slope \(\frac{5}{2}\) the "negative reciprocal"

OpenStudy (whpalmer4):

Ah, the \ ( vs. \ [ was the trick, got it, thanks! If the slope of the first line is a fraction, invert it and change the sign.

OpenStudy (anonymous):

yeah in line latex use \( instead of \[ works in documents too

OpenStudy (anonymous):

that is, you can use \( instead of $

OpenStudy (whpalmer4):

Too bad they don't support the even shorter $eqn$

Parth (parthkohli):

But they do support `$$ $$`. $$\text{Hi.}$$

OpenStudy (anonymous):

if i remember correctly (and i could easily be wrong) $ is a hack

Parth (parthkohli):

No, $ is perfect.

Parth (parthkohli):

And original.

OpenStudy (anonymous):

like i said, i could easily be wrong

OpenStudy (whpalmer4):

Yep, $ is a long-time component of the math environment. I haven't written any TeX or LaTeX documents since the days when Knuth was still occasionally working on TeX, MetaFont, Tangle, etc. Thanks for the info!

OpenStudy (anonymous):

damn you must be as old as i am !

OpenStudy (laddiusmaximus):

im not seeing it.

OpenStudy (whpalmer4):

What aren't you seeing? You understand how to find the slope of the original line, right? Two perpendicular lines will have a product of their slopes = -1. Think of the graphs of \(y=x\) and \(y = -x\) — they make a nice big symmetrical X over the origin, and are perpendicular. The slope of \(y=x\) is 1, and the slope of \(y = -x\) is -1, because they are in slope-intercept form:$$y=mx+b$$where m is slope and b is y-intercept. Now, if you have two numbers \(a,b\) where \(a*b = -1\), you can find b from a by \[b = -\frac{1}{a}\] Dividing by a fraction is the same as multiplying by the reciprocal of the fraction, so \[-\frac{1}{-\frac{2}{5}} = -1 * -\frac{5}{2} = \frac{5}{2}\]